A survey asks a random sample of 1500 adults in Ohio if they support an increase in the state sales tax from 5% to 6%, with the additional revenue going to education. Let ^ p denote the proportion in the sample who say they support the increase. Suppose that 36% of all adults in Ohio support the increase. The standard deviation of the sampling distribution is . Round your answer to four decimal places.
Answer #1

Solution

Given that,

p = 0.36

1 - p = 1 - 0.36=0.64

n = 1500

= p =0.36

=  [p ( 1 - p ) / n] =   [(0.36*0.64) /1500 ] = 0.0124

answer = 0.0124